Monthly food expenditures for families in rodriguez rizal averaged php6,500 with




1
Expert's answer
2022-02-14T13:02:48-0500

"n=7\\\bar x={\sum x\over n}={36\over7}=5.14285714\\ s^2={\sum x^2-{(\sum(x))^2\over n}\over n-1}={220-185.142857\over6}=5.8095\\s=\sqrt{s^2}=\sqrt{5.8095}=2.4103"

The hypotheses are,

"H_0:\mu=4.25\\vs\\H_1:\mu\gt4.25"

The test statistic is,

"t={\bar x-\mu\over {s\over\sqrt{n}}}={5.14-4.25\over{2.4103\over\sqrt{7}}}={0.8929\over0.911}=0.98"

The critical value is,

"t_{\alpha,n-1}=t_{0.1,6}=1.439756"

The null hypothesis is rejected if, "t\gt t_{0.1,6}"

Since "t=0.98\lt t_{0.1,6}=1.439756", the null hypothesis is not rejected. Therefore, we can conclude that there is no sufficient evidence to show that the average family size is more than the national average at 1% level of significance.